# Efficient Point Cloud Classification via Offline Distillation Framework and Negative-Weight Self-Distillation Technique

### v1

@@ 4.4 Effects of Distillation Weights on Model Performance:
4.5 Impact of Distillation Techniques on Feature Space Representation
@@
In this section, we present a t-distributed Stochastic Neighbor Embedding (t-SNE) visualization analysis of four models for point cloud classification. The models include: (a) the original teacher model, PointViG [25]; (b) PointViG-Distil, a compressed version without any distillation techniques; (c) PointViG-Distil with only teacher-student distillation; and (d) the full PointViG-Distil model incorporating both teacher-student distillation and negative-weight self-distillation. Each model’s encoded features are visualized in separate plots, labeled (a), (b), (c), and (d) in Fig. 2, respectively. Ten representative regions were identified and consistently labeled across all visualizations to facilitate a direct comparison of class separability and feature representation.
@@
Figure 2: t-SNE visualization of encoded features for (a) PointViG [25] (teacher model), (b) PointViG-Distil (no distillation), (c) PointViG-Distil (teacher-student distillation only), and (d) PointViG-Distil (teacher-student and negative-weight self-distillation). Ten representative regions are consistently labeled across all plots for comparison of class separability and feature representation. This figure is best viewed in an enlarged format for clarity.
The visualizations revealed distinct patterns in the feature space for each model. Model-a, the teacher model PointViG [25], exhibited well-dispersed clusters with clear boundaries, indicating high separability between categories. In contrast, Model-b, the compressed version without distillation, showed less distinct cluster boundaries, suggesting reduced separability compared to the teacher model. Model-c, which introduced teacher-student distillation to Model-b, closely resembled Model-a, particularly for clusters 1, 2, 3, 4, 6, 7, 8, and 10, indicating that the student model effectively emulated the teacher’s feature representation.
Model-d, the full PointViG-Distil model with both distillation techniques, showed an expanded distribution range for all clusters compared to Models a and c. Notably, smaller clusters, which were centrally located, were more dispersed in Model-d, suggesting improved separability. Additionally, in Model-d, clusters 1, 2, 3, 5, 7, and 10 displayed a substantial reduction in category overlap, with the majority of instances where overlap occurs being confined to the periphery of the clusters, thereby indicating a mitigated inter-category interference. The observed improvements in Model-d can be attributed to the combined effects of teacher-student distillation and negative-weight self-distillation. The teacher-student distillation allows the student model to inherit the teacher’s knowledge, enhancing feature representation and class separability. The negative-weight self-distillation further refines the model by encouraging diversity in the feature space, reducing the likelihood of converging on suboptimal solutions and promoting generalization.
The t-SNE visualization analysis underscores the efficacy of our distillation strategies in enhancing the PointViG-Distil model’s feature encoding capabilities. The incremental integration of distillation techniques, from Model-b to Model-d, has led to a progressive enhancement in class separability and feature representation. The full PointViG-Distil model, with its combined distillation approaches, offers a compelling balance between model complexity and performance, making it well-suited for applications with limited computational resources.

### v2

@@ 4.4 Effects of Distillation Weights on Model Performance:
4.5 Visualization Analysis of Encoder Features
@@
In this section, we present a t-distributed Stochastic Neighbor Embedding (t-SNE) visualization analysis of four models for point cloud classification. The models include: (a) the original teacher model, PointViG [25]; (b) PointViG-Distil, a compressed version without any distillation techniques; (c) PointViG-Distil with only teacher-student distillation; and (d) the full PointViG-Distil model incorporating both teacher-student distillation and negative-weight self-distillation. Each model’s encoded features are visualized in separate plots, labeled (a), (b), (c), and (d) in Fig. 2, respectively.
@@
Figure 2: t-SNE visualization of encoded features for (a) PointViG [25] (teacher model), (b) PointViG-Distil (no distillation), (c) PointViG-Distil (teacher-student distillation only), and (d) PointViG-Distil (teacher-student and negative-weight self-distillation). This figure is best viewed in an enlarged format for clarity.
In Fig. 2(a), the PointViG teacher model shows well-defined clusters, reflecting strong intra-cluster aggregation and distinct class separability, setting the benchmark for comparison. Moving to Fig. 2(b), where no distillation operations are applied, the student model exhibits reduced clustering efficiency. Several clusters are less compact, and some have more stray points, indicating weaker feature representations and reduced class separability.
Fig. 2(c), representing the student model with only teacher-student distillation, shows significant improvement over Model-b, closely mimicking the clustering behaviour of the teacher model. Compared to Model-b, the clusters of Model-c exhibit tighter formations and clearer separations, demonstrating that the distillation process helps the student model replicate the teacher’s feature distribution, enhancing its representational power.
Fig. 2(d) depicts the full PointViG-Distil model, incorporating both teacher-student and self-distillation. Although the accuracy of this model is marginally lower than the teacher model, the t-SNE visualisation reveals more balanced cluster distributions. Compared to Model-c, the clusters of PointViG-Distil exhibit fewer outliers and clearer boundaries, reducing overlap between classes. This improvement in cluster separability and reduction in free points confirms the efficacy of negative-weight self-distillation in improving the generalization capacity of the student model.
In summary, the comparison across these four models highlights the progression in feature distribution quality with the addition of distillation techniques, validating the role of both teacher-student and negative-weight self-distillation in improving the robustness and generalization of PointViG-Distil.
4.6 Visualization Analysis of Logit Outputs
In this section, we present a t-SNE visualization analysis of the logit outputs for four models in the context of point cloud classification. The models include: (a) the original teacher model, PointViG [25]; (b) PointViG-Distil, a compressed version without any distillation techniques; (c) PointViG-Distil with only teacher-student distillation; and (d) the full PointViG-Distil model incorporating both teacher-student distillation and negative-weight self-distillation. Each model’s logits are visualized in separate plots, labeled (a), (b), (c), and (d) in Fig. 3, respectively. Ten representative clusters were identified and consistently labeled across all visualizations to facilitate a direct comparison of the models’ confidence levels and decision boundaries.
Figure 3: t-SNE visualization of encoded features for (a) PointViG [25] (teacher model), (b) PointViG-Distil (no distillation), (c) PointViG-Distil (teacher-student distillation only), and (d) PointViG-Distil (teacher-student and negative-weight self-distillation). Ten representative regions are consistently labeled across all plots for comparison of confidence levels and decision boundaries. This figure is best viewed in an enlarged format for clarity.
The visualizations revealed clear differences in the confidence and decision boundary delineation across models. Model-a, the teacher model PointViG [25], showed compact and well-delineated clusters. This tight grouping reflects high classification confidence, as the decision boundaries between classes are sharp, and the logits for each class are confidently grouped. The clear boundaries indicate that the teacher model makes strong, well-defined decisions, leaving little ambiguity between classes.
In contrast, Model-b, the compressed version without any distillation techniques, exhibited more dispersed clusters with fuzzier boundaries. This dispersion indicates a drop in classification confidence and less sharply defined decision boundaries. The logits from this model show increased overlap between categories, suggesting that the compressed model without distillation struggles to maintain confident, unambiguous decision-making between categories, leading to more uncertainty near the decision boundaries.
Model-c, which reintroduces teacher-student distillation, shows significant improvement in confidence and decision-making clarity. The clusters, particularly 1, 2, 3, 4, 6, 7, and 9, more closely resemble those from Model-a. This indicates that the addition of teacher-student distillation helps the student model mimic the teacher’s decision boundaries and confidence levels effectively. The model recaptures much of the teacher’s ability to make confident and well-separated decisions, reducing overlap between different classes and enhancing decision reliability.
Model-d, the full PointViG-Distil model incorporating both teacher-student and self-distillation, further refines the logit space. The overall distribution of clusters is more spread out compared to Model-c. This spread indicates that the model allows for more flexibility in its decision boundaries, while maintaining high confidence in its classifications. Importantly, the overlap of points between different categories is minimized, particularly in critical areas like clusters 1, 2, 5, 7, and 10. This suggests that the model achieves a balance between strong classification confidence and broader decision regions, allowing for improved handling of difficult-to-classify samples compared to Model-c.
The t-SNE visualization analysis of the logit outputs highlight the progressive improvements from removing distillation (Model-b) to adding teacher-student distillation (Model-c) and ultimately introducing negative-weight self-distillation (Model-d), each contributing to the model’s ability to make confident and precise decisions.